# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

### v1

@@ Abstract
We derive a new Carleman estimate for a general backward stochastic parabolic equation with dynamic boundary conditions, incorporating weak divergence source terms in both the bulk and surface equations. This estimate is obtained through two main steps: first, by refining a known Carleman estimate for backward stochastic parabolic equations to explicitly account for the dependence of the parameters on the final control time $T$; second, by applying a duality technique to address weak divergence source terms. As an application, we prove the null and approximate controllability of forward stochastic parabolic equations with dynamic boundary conditions, which involve both reaction and convection terms with bounded adapted coefficients, as well as general second-order parabolic operators. Additionally, we provide an explicit estimate for the null-controllability cost in terms of the final control time $T$ and the coefficients of the equation.
Cadi Ayyad University, National School of Applied Sciences, Laboratory of Mathematics, Modeling and Automatic Systems, B.P. 575, Marrakesh, Morocco. Email: s.boulite@uca.ma
Dipartimento di Matematica, Università degli Studi di Salerno, Via Giovanni Paolo II, 132, 84084 Fisciano (SA), Italy. Email: aelgrou@unisa.it,  arhandi@unisa.it
§Cadi Ayyad University, Faculty of Sciences Semlalia, Laboratory of Mathematics, Modeling and Automatic Systems, B.P. 2390, Marrakesh, Morocco. Email: maniar@uca.ma
University Mohammed VI Polytechnic, Vanguard Center, Benguerir, Morocco.
Email: Lahcen.Maniar@um6p.ma
Keywords: Carleman estimates; Stochastic parabolic equations; Null controllability; Approximate controllability; Observability inequality; Unique continuation property; Dynamic boundary conditions.
@@ 1 Introduction and Main Results
Let $T>0$, $G\subset\mathbb{R}^{N}$ be a given nonempty open bounded domain with a smooth boundary $\Gamma=\partial G$, and $N\geq 2$. Let $G_{0}\Subset G$ (i.e., $\overline{G_{0}}\subset G$, where $\overline{G_{0}}$ denotes the closure of $G_{0}$) be a nonempty open subset strictly contained in $G$, with $\mathbbm{1}_{G_{0}}$ denoting the characteristic function of $G_{0}$. Define the following sets:
$Q=(0,T)\times G,\quad\Sigma=(0,T)\times\Gamma,\quad\text{and}\quad Q_{0}=(0,T)\times G_{0}.$
Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a fixed complete filtered probability space on which a one-dimensional standard Brownian motion $W(\cdot)$ is defined, such that $\{\mathcal{F}_{t}\}_{t\geq 0}$ is the natural filtration generated by $W(\cdot)$ and augmented by all the $\mathbb{P}$-null sets in $\mathcal{F}$. Let $H$ be a Banach space, and let $C([0,T];H)$ be the Banach space of all $H$-valued continuous functions defined on $[0,T]$. For some sub-sigma algebra $\mathcal{G}\subset\mathcal{F}$, we denote by $L^{2}_{\mathcal{G}}(\Omega;H)$ the Banach space of all $H$-valued $\mathcal{G}$-measurable random variables $X$ such that $\mathbb{E}\left[\|X\|_{H}^{2}\right]<\infty$, with the canonical norm. We also denote by $L^{2}_{\mathcal{F}}(0,T;H)$ the Banach space consisting of all $H$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted processes $X(\cdot)$ such that $\mathbb{E}\left[\|X(\cdot)\|^{2}_{L^{2}(0,T;H)}\right]<\infty$, with the canonical norm. Moreover, we denote by $L^{\infty}_{\mathcal{F}}(0,T;H)$ the Banach space consisting of all $H$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted essentially bounded processes. Finally, we denote by $L^{2}_{\mathcal{F}}(\Omega;C([0,T];H))$ the Banach space consisting of all $H$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted continuous processes $X(\cdot)$ such that $\mathbb{E}\left[\|X\|^{2}_{C([0,T];H)}\right]<\infty$, with the canonical norm.
Consider the following Hilbert space
@@
where $dx$ denotes the Lebesgue measure on $G$ and $d\sigma$ denotes the surface measure on $\Gamma$, equipped with the following inner product:
$\langle(y,y_{\Gamma}),(z,z_{\Gamma})\rangle_{\mathbb{L}^{2}}=\langle y,z\rangle_{L^{2}(G)}+\langle y_{\Gamma},z_{\Gamma}\rangle_{L^{2}(\Gamma)}.$
We assume that the diffusion matrices $\mathcal{A}=(a^{jk})_{j,k=1,2,\cdots,N}$ and $\mathcal{A}_{\Gamma}=(b^{jk})_{j,k=1,2,\cdots,N}$ satisfy the following assumptions:
@@
    $\mathcal{A}\in L^{\infty}_{\mathcal{F}}(\Omega;C^{1}([0,T];W^{2,\infty}(\overline{G};\mathbb{R}^{N\times N})))$, and $a^{jk}=a^{kj}$ for all $j,k=1,2,\cdots,N$.
@@
    $\mathcal{A}_{\Gamma}\in L^{\infty}_{\mathcal{F}}(\Omega;C^{1}([0,T];W^{1,\infty}(\Gamma;\mathbb{R}^{N\times N})))$, and $b^{jk}=b^{kj}$ for all $j,k=1,2,\cdots,N$.
@@
    $\begin{aligned} \langle\mathcal{A}(\omega,t,x)\xi,\xi\rangle_{\mathbb{R}^{N}}&\geq\beta_{0}|\xi|^{2}\quad\text{for any}\quad(\omega,t,x,\xi)\in\Omega\times(0,T)\times\overline{G}\times\mathbb{R}^{N},\\ \langle\mathcal{A}_{\Gamma}(\omega,t,x)\xi,\xi\rangle_{\mathbb{R}^{N}}&\geq\beta_{0}|\xi|^{2}\quad\text{for any}\quad(\omega,t,x,\xi)\in\Omega\times(0,T)\times\Gamma\times\mathbb{R}^{N},\end{aligned}$ (1.1)
    where $\langle\cdot,\cdot\rangle_{\mathbb{R}^{N}}$ (resp. $|\cdot|$) is the Euclidean inner product (resp. norm) in $\mathbb{R}^{N}$.
Throughout this paper, $C$ denotes a positive constant depending only on $G$, $G_{0}$, $\mathcal{A}$, and $\mathcal{A}_{\Gamma}$, which may change from one place to another.
Consider the following non-homogeneous linear backward stochastic parabolic equation subject to dynamic boundary conditions
$\begin{cases}\begin{array}[]{ll}dz+\nabla\cdot(\mathcal{A}\nabla z)\,dt=(F_{1}+\nabla\cdot F)\,dt+Z\,dW(t)&\text{in}\,Q,\\ dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt=(F_{2}-F\cdot\nu+\nabla_{\Gamma}\cdot F_{\Gamma})\,dt+\widehat{Z}\,dW(t)&\text{on}\,\Sigma,\\ z_{\Gamma}=z|_{\Gamma}&\text{on}\,\Sigma,\\ (z,z_{\Gamma})|_{t=T}=(z_{T},z_{\Gamma,T})&\text{in}\,G\times\Gamma,\end{array}\end{cases}$ (1.2)
@@
$\partial^{\mathcal{A}}_{\nu}z=(\mathcal{A}\nabla z\cdot\nu)|_{\Sigma}=\sum_{j,k=1}^{N}a^{jk}\frac{\partial z}{\partial x_{j}}\nu^{k},$
@@
represents the conormal derivative of $z$ with respect to $\mathcal{A}$, where $\nu=(\nu^{1},\cdots,\nu^{N})$ is the outer unit normal vector at $\Gamma$. This conormal derivative acts as the coupling term between the bulk equation in $Q$ and the surface equation on $\Sigma$, meaning the bulk equation is controlled directly, while the surface equation is governed through the conormal derivative $\partial_{\nu}^{\mathcal{A}}z$.
The operators $\nabla\cdot$ (in $G$) and $\nabla_{\Gamma}\cdot$ (on $\Gamma$) are the divergence operators. If $\mathcal{A}=I_{N}$, the identity matrix, then the conormal derivative $\partial_{\nu}^{\mathcal{A}}z$ simplifies to the standard normal derivative $\partial_{\nu}z$, and the bulk diffusion term $\nabla\cdot(\mathcal{A}\nabla z)$ becomes the Laplacian $\Delta z$. The surface term $\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})$ captures surface diffusion effects on $\Gamma$, where the tangential gradient $\nabla_{\Gamma}z_{\Gamma}:=\nabla z-(\partial_{\nu}z)\nu$ is used. In this case, if $\mathcal{A}_{\Gamma}=I_{N}$, the surface diffusion term simplifies to the Laplace-Beltrami operator $\Delta_{\Gamma}z_{\Gamma}$ on $\Gamma$. For further details on these geometric tools and on dynamic boundary conditions, we refer the reader to [17, 20, 28, 33].
Similarly to the standard Sobolev spaces $H^{k}(G)$ ($k=1,2$), we introduce the Sobolev spaces $H^{k}(\Gamma)$ ($k=1,2$) on the boundary $\Gamma$ as follows:
@@
$H^{k}(\Gamma)=\left\{z_{\Gamma}\in L^{2}(\Gamma)\mid\nabla_{\Gamma}z_{\Gamma}\in H^{k-1}(\Gamma;\mathbb{R}^{N})\right\},\quad k=1,2,$
which are endowed with the following norms:
@@
$|z_{\Gamma}|_{H^{1}(\Gamma)}=\sqrt{\langle z_{\Gamma},z_{\Gamma}\rangle_{H^{1}(\Gamma)}},\quad\text{where}\quad\langle y_{\Gamma},z_{\Gamma}\rangle_{H^{1}(\Gamma)}=\int_{\Gamma}y_{\Gamma}z_{\Gamma}\,d\sigma+\int_{\Gamma}\nabla_{\Gamma}y_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma,$
@@
$|z_{\Gamma}|_{H^{2}(\Gamma)}=\sqrt{\langle z_{\Gamma},z_{\Gamma}\rangle_{H^{2}(\Gamma)}},\quad\text{where}\quad\langle y_{\Gamma},z_{\Gamma}\rangle_{H^{2}(\Gamma)}=\int_{\Gamma}y_{\Gamma}z_{\Gamma}\,d\sigma+\int_{\Gamma}\Delta_{\Gamma}y_{\Gamma}\,\Delta_{\Gamma}z_{\Gamma}\,d\sigma.$
In the sequel, we also define the following Hilbert spaces:
$\mathbb{H}^{k}=\{(z,z_{\Gamma})\in H^{k}(G)\times H^{k}(\Gamma):z_{\Gamma}=z|_{\Gamma}\},\quad k=1,2,$
equipped with the usual inner product
$\langle(y,y_{\Gamma}),(z,z_{\Gamma})\rangle_{\mathbb{H}^{k}}=\langle y,z\rangle_{H^{k}(G)}+\langle y_{\Gamma},z_{\Gamma}\rangle_{H^{k}(\Gamma)}.$
In this paper, the following property
$(\textbf{z},\textbf{z}_{\Gamma})=(z_{1},z_{2})\quad\text{in}\;G\times\Gamma,\quad\textnormal{a.s.},$
is understood as follows:
$\textbf{z}=z_{1}\quad\text{in}\;G\quad\text{and}\quad\textbf{z}_{\Gamma}=z_{2}\quad\text{on}\;\Gamma,\quad\textnormal{a.s.}$
Remark 1.1.
@@
In this paper, the concept of vectors on the boundary $\Gamma$ is understood in the context of vector fields on the manifold $\Gamma$. More precisely, these vectors are functions $\gamma:\Gamma\to\mathbb{R}^{N}$ such that for each $x\in\Gamma$, the vector $\gamma(x)$ lies in the tangent space $T_{x}\Gamma$ at $x$, defined by
$T_{x}\Gamma=\{p\in\mathbb{R}^{N}\mid p\cdot\nu(x)=0\}.$
This assumption allows for simplifications in the computation of certain boundary terms via the tangential divergence formula:
$\int_{\Gamma}(\nabla_{\Gamma}\cdot\gamma)\,z_{\Gamma}\,d\sigma=-\int_{\Gamma}\gamma\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma,$
for $\gamma\in H^{1}(\Gamma;\mathbb{R}^{N})$ and $z_{\Gamma}\in H^{1}(\Gamma)$. However, when the vector field $\gamma:\Gamma\to\mathbb{R}^{N}$ is not necessarily tangential, one needs to use the generalized divergence formula (see [16, Proposition 5.4.9]):
$\int_{\Gamma}(\nabla_{\Gamma}\cdot\gamma)\,z_{\Gamma}\,d\sigma=-\int_{\Gamma}\gamma\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma+\int_{\Gamma}(\nabla_{\Gamma}\cdot\nu)\,(\gamma\cdot\nu)\,z_{\Gamma}\,d\sigma.$
This identity introduces additional boundary-like terms that need to be carefully treated.
@@
By the duality method as in [4, Section 2], it can be shown that the equation (1.2) is well-posed. Specifically, for all given data $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, $F_{1}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$, $F\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G;\mathbb{R}^{N}))$, $F_{2}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, and $F_{\Gamma}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma;\mathbb{R}^{N}))$, the system (1.2) has a unique weak solution
@@
$(z,z_{\Gamma};Z,\widehat{Z})\in\left(L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})\right)\times L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2}).$

### v2

# Carleman Estimates for Backward Anisotropic Stochastic Parabolic Equations with General Dynamic Boundary Conditions and Applications

@@ Abstract
We investigate a backward anisotropic stochastic parabolic equation with general dynamic boundary conditions, where the drift involves both $\mathbb{L}^{2}$ and $\mathbb{H}^{-1}$ bulk–surface terms. We first establish the well-posedness of this equation. Subsequently, we derive a new Carleman estimate through a two-step approach. In the first step, using a weighted identity method together with a careful treatment of the boundary integral terms arising from the dynamic boundary conditions, we obtain an intermediate Carleman estimate for backward anisotropic stochastic parabolic equations without weak divergence source terms. In the second step, a duality method combined with suitable optimization techniques is employed to incorporate the weak divergence source terms. As applications of the derived Carleman estimate, we address two control problems. First, we establish null controllability for forward anisotropic stochastic parabolic equations with general dynamic boundary conditions. These equations involve both reaction and convection terms, with adapted, bounded stochastic bulk–surface coefficients. Moreover, we provide an explicit estimate of the null controllability cost, i.e., a bound on the minimal norm of controls required to drive the system to zero at the terminal time $T$. Second, we study an insensitizing control problem for this class of equations. The goal is to determine controls for systems with partially unknown initial data such that a given energy functional remains insensitive to small perturbations of these data. In this work, the functional involves the norm of the state over a localized bulk–surface region, together with the norm of its tangential gradient over a localized boundary region.
Keywords: Carleman estimates; Anisotropic stochastic parabolic equations; Controllability; Dynamic boundary conditions; Surface diffusion; Control costs; Insensitizing control.
@@ 1 Introduction and Main Results
Fix a final time $T>0$, and let $G\subset\mathbb{R}^{N}$ be a nonempty bounded open domain with boundary $\Gamma=\partial G$ of class $C^{4}$, where $N\geq 2$. Let $G_{0},\mathcal{O}\Subset G$ be nonempty open subsets strictly contained in $G$, and let $\mathcal{O}^{1}_{\Gamma}$ and $\mathcal{O}^{2}_{\Gamma}$ be any nonempty open subsets of $\Gamma$. Denote by $\mathbbm{1}_{S}$ the characteristic function of a set $S\subset\overline{G}$, and by $\mathbb{E}[\cdot]$ the expectation operator with respect to the underlying probability space. The following time-space domains are considered:
$Q:=(0,T)\times G,\qquad\Sigma:=(0,T)\times\Gamma,\qquad Q_{0}:=(0,T)\times G_{0}.$
Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a fixed complete filtered probability space on which a one-dimensional standard Brownian motion $W(\cdot)$ is defined, such that $\{\mathcal{F}_{t}\}_{t\geq 0}$ is the natural filtration generated by $W(\cdot)$ and augmented by all the $\mathbb{P}$-null sets in $\mathcal{F}$. Let $\mathcal{X}$ be a Banach space, and let $C([0,T];\mathcal{X})$ be the Banach space of all $\mathcal{X}$-valued continuous functions defined on $[0,T]$. For some sub-sigma algebra $\mathcal{G}\subset\mathcal{F}$, we denote by $L^{2}_{\mathcal{G}}(\Omega;\mathcal{X})$ the Banach space of all $\mathcal{X}$-valued $\mathcal{G}$-measurable random variables $X$ such that $\mathbb{E}\left[\|X\|_{\mathcal{X}}^{2}\right]<\infty$, with the canonical norm. We also denote by $L^{2}_{\mathcal{F}}(0,T;\mathcal{X})$ the Banach space consisting of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted processes $X(\cdot)$ such that $\mathbb{E}\left[\int_{0}^{T}\|X(t)\|^{2}_{\mathcal{X}}\,dt\right]<\infty$, with the canonical norm. Moreover, we denote by $L^{\infty}_{\mathcal{F}}(0,T;\mathcal{X})$ the Banach space consisting of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted essentially bounded processes. Finally, we denote by $L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathcal{X}))$ the Banach space consisting of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted continuous processes $X(\cdot)$ such that $\mathbb{E}\left[\max_{0\leq t\leq T}\|X(t)\|^{2}_{\mathcal{X}}\right]<\infty$, with the canonical norm.
Consider the following product space:
@@
where $dx$ denotes the Lebesgue measure on $G$ and $d\sigma$ denotes the surface measure on $\Gamma$. One can easily verify that $\mathbb{L}^{2}$ is a Hilbert space with the usual inner product
$\langle(y,y_{\Gamma}),(z,z_{\Gamma})\rangle_{\mathbb{L}^{2}}:=\langle y,z\rangle_{L^{2}(G)}+\langle y_{\Gamma},z_{\Gamma}\rangle_{L^{2}(\Gamma)}.$
We assume that the functions $a^{jk}:\overline{G}\to\mathbb{R}$ and $a_{\Gamma}^{jk}:\Gamma\to\mathbb{R}$ ($j,k=1,2,\cdots,N$), which represent the diffusion coefficients in the domain and on the boundary, respectively, satisfy the following assumptions:
@@
    $a^{jk}\in C^{2}(\overline{G})$, and $a^{jk}=a^{kj}$ for all $j,k=1,2,\cdots,N$.
@@
    $a_{\Gamma}^{jk}\in C^{1}(\Gamma)$, and $a_{\Gamma}^{jk}=a_{\Gamma}^{kj}$ for all $j,k=1,2,\cdots,N$.
@@
    $\begin{aligned} \sum_{j,k=1}^{N}a^{jk}(x)\eta_{j}\eta_{k}&\geq\beta_{0}|\eta|^{2}\quad\textnormal{for all }(x,\eta)\equiv(x_{1},\cdots,x_{N},\eta_{1},\cdots,\eta_{N})\in\overline{G}\times\mathbb{R}^{N},\\ \sum_{j,k=1}^{N}a_{\Gamma}^{jk}(x)\eta_{j}\eta_{k}&\geq\beta_{0}|\eta|^{2}\quad\textnormal{for all }(x,\eta)\equiv(x_{1},\cdots,x_{N},\eta_{1},\cdots,\eta_{N})\in\Gamma\times\mathbb{R}^{N},\end{aligned}$ (1.1)
    where $|\cdot|$ stands for the Euclidean norm in $\mathbb{R}^{N}$.
Throughout this paper, $C$ denotes a generic positive constant depending only on $G$, $G_{0}$, $\beta_{0}$, $M$, and $M_{\Gamma}$, which may change from line to line, with
$M=\max_{1\leq j,k\leq N}\|a^{jk}\|_{C^{2}(\overline{G})},\qquad M_{\Gamma}=\max_{1\leq j,k\leq N}\|a_{\Gamma}^{jk}\|_{C^{1}(\Gamma)}.$
The main goal of this paper is to establish the well-posedness and to derive a new Carleman estimate for the following nonhomogeneous linear backward stochastic parabolic equation subject to general dynamic boundary conditions:
$\begin{cases}dz+\displaystyle\sum_{j,k=1}^{N}\frac{\partial}{\partial x_{j}}\Big(a^{jk}(x)\frac{\partial z}{\partial x_{k}}\Big)\,dt=\Big(F_{1}+\sum_{j=1}^{N}\frac{\partial F_{j}}{\partial x_{j}}\Big)\,dt+Z\,dW(t),&\textnormal{in }Q,\\[2.84526pt] dz_{\Gamma}+\displaystyle\sum_{j,k=1}^{N}D_{j}\Big(a_{\Gamma}^{jk}(x)D_{k}z_{\Gamma}\Big)\,dt-\displaystyle\sum_{j,k=1}^{N}a^{jk}(x)\frac{\partial z}{\partial x_{j}}\nu^{k}\,dt=\Big(F_{2}-\displaystyle\sum_{j=1}^{N}F_{j}\nu^{j}+\sum_{j=1}^{N}D_{j}{F_{\Gamma}}_{j}\Big)\,dt+\widehat{Z}\,dW(t),&\textnormal{on }\Sigma,\\[2.84526pt] z_{\Gamma}=z|_{\Gamma},&\textnormal{on }\Sigma,\\[2.84526pt] (z,z_{\Gamma})|_{t=T}=(z_{T},z_{\Gamma,T}),&\textnormal{in }G\times\Gamma.\end{cases}$ (1.2)
@@
$F_{1}$ $\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G)),\qquad F_{2}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)),$
$F_{j}$ $\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G)),\qquad{F_{\Gamma}}_{j}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)),\qquad j=1,2,\cdots,N.$
In (1.2), $\frac{\partial}{\partial x_{j}}$ and $D_{j}$ denote, respectively, the standard partial derivative with respect to $x_{j}$ and the $j$-th tangential derivative operator on $\Gamma$, defined by
$D_{j}:=\frac{\partial}{\partial x_{j}}-\nu^{j}\sum_{k=1}^{N}\nu^{k}\frac{\partial}{\partial x_{k}},\qquad j=1,2,\cdots,N,$
where $\nu^{j}$ is the $j$-th component of the unit outward normal vector $\nu=(\nu^{1},\dots,\nu^{N})$ to the boundary $\Gamma$. The tangential derivative represents the part of the standard partial derivative that is tangent to the surface, obtained by subtracting the normal component from the standard derivative. For further details on tangential derivative operators, we refer the reader to [12]. Throughout this work, the boundary $\Gamma$ of $G$ is regarded as a (N1)(N-1)-dimensional compact Riemannian submanifold induced by the natural embedding $\Gamma\hookrightarrow\mathbb{R}^{N}$, and the associated surface measure $d\sigma$.
It is straightforward to see that, by introducing the notations
$\mathcal{A}=(a^{jk})_{1\leq j,k\leq N},\qquad\mathcal{A}_{\Gamma}=(a_{\Gamma}^{jk})_{1\leq j,k\leq N},\qquad F=(F_{j})_{1\leq j\leq N},\qquad F_{\Gamma}=({F_{\Gamma}}_{j})_{1\leq j\leq N},$
the system (1.2) can be equivalently rewritten in the following compact form:
$\begin{cases}\begin{array}[]{ll}dz+\nabla\cdot(\mathcal{A}(x)\nabla z)\,dt=(F_{1}+\nabla\cdot F)\,dt+Z\,dW(t)&\textnormal{in }Q,\\[3.00003pt] dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}(x)\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt=(F_{2}-F\cdot\nu+\nabla_{\Gamma}\cdot F_{\Gamma})\,dt+\widehat{Z}\,dW(t)&\textnormal{on }\Sigma,\\[3.00003pt] z_{\Gamma}=z|_{\Gamma}&\textnormal{on }\Sigma,\\[3.00003pt] (z,z_{\Gamma})|_{t=T}=(z_{T},z_{\Gamma,T})&\textnormal{in }G\times\Gamma,\end{array}\end{cases}$ (1.3)
where the boundary term
$\partial_{\nu}^{\mathcal{A}}z:=(\mathcal{A}\nabla z\cdot\nu)|_{\Sigma}=\sum_{j,k=1}^{N}a^{jk}\frac{\partial z}{\partial x_{j}}\nu^{k}$
@@
represents the conormal derivative of $z$ with respect to $\mathcal{A}$. This conormal derivative serves as a coupling term between the bulk equation in $Q$ and the surface equation on $\Sigma$, so that the bulk equation is governed directly, while the surface equation is influenced through $\partial_{\nu}^{\mathcal{A}}z$. The operators $\nabla\cdot F$ (in $G$) and $\nabla_{\Gamma}\cdot F_{\Gamma}$ (on $\Gamma$) denote, respectively, the standard divergence in $G$ and the tangential divergence on $\Gamma$, defined by
@@
$\nabla\cdot F:=\sum_{j=1}^{N}\frac{\partial F_{j}}{\partial x_{j}},\qquad\nabla_{\Gamma}\cdot F_{\Gamma}:=\sum_{j=1}^{N}D_{j}{F_{\Gamma}}_{j}=\sum_{j=1}^{N}\left(\frac{\partial{F_{\Gamma}}_{j}}{\partial x_{j}}-\nu^{j}\sum_{k=1}^{N}\nu^{k}\frac{\partial{F_{\Gamma}}_{j}}{\partial x_{k}}\right).$
The surface term $\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})$ models diffusion on $\Gamma$, where the tangential gradient is defined by
$\nabla_{\Gamma}z_{\Gamma}:=(D_{j}z_{\Gamma})_{1\leq j\leq N}=\nabla z-(\partial_{\nu}z)\nu,$
which is the projection of the gradient $\nabla z$ onto the tangent plane at $x\in\Gamma$. For further details on these geometric tools and on dynamic boundary conditions, we refer the reader to [21, 25, 36, 41].
Similarly to the standard Sobolev spaces $H^{k}(G)$ ($k=1,2$), we introduce the Sobolev spaces $H^{k}(\Gamma)$ ($k=1,2$) on the boundary $\Gamma$ by
$H^{k}(\Gamma)=\bigl\{z_{\Gamma}\in L^{2}(\Gamma)\,\big|\,\nabla_{\Gamma}z_{\Gamma}\in H^{k-1}(\Gamma;\mathbb{R}^{N})\bigr\},\qquad k=1,2.$
These spaces are endowed with the following norms:
@@
$\|z_{\Gamma}\|_{H^{1}(\Gamma)}:=\langle z_{\Gamma},z_{\Gamma}\rangle_{H^{1}(\Gamma)}^{1/2},\quad\textnormal{where}\quad\langle y_{\Gamma},z_{\Gamma}\rangle_{H^{1}(\Gamma)}=\int_{\Gamma}y_{\Gamma}z_{\Gamma}\,d\sigma+\int_{\Gamma}\nabla_{\Gamma}y_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma,$
@@
$\|z_{\Gamma}\|_{H^{2}(\Gamma)}:=\langle z_{\Gamma},z_{\Gamma}\rangle_{H^{2}(\Gamma)}^{1/2},\quad\textnormal{where}\quad\langle y_{\Gamma},z_{\Gamma}\rangle_{H^{2}(\Gamma)}=\int_{\Gamma}y_{\Gamma}z_{\Gamma}\,d\sigma+\int_{\Gamma}\Delta_{\Gamma}y_{\Gamma}\,\Delta_{\Gamma}z_{\Gamma}\,d\sigma.$
In the sequel, we introduce the following Hilbert spaces:
$\mathbb{H}^{k}:=\bigl\{(z,z_{\Gamma})\in H^{k}(G)\times H^{k}(\Gamma)\,\big|\,z_{\Gamma}=z|_{\Gamma}\bigr\},\qquad k=1,2,$
which are endowed with the natural inner product
$\langle(y,y_{\Gamma}),(z,z_{\Gamma})\rangle_{\mathbb{H}^{k}}:=\langle y,z\rangle_{H^{k}(G)}+\langle y_{\Gamma},z_{\Gamma}\rangle_{H^{k}(\Gamma)}.$
Moreover, throughout this paper we denote by $(H^{k}(G))^{\prime}$, $H^{-k}(\Gamma)$, and $\mathbb{H}^{-k}$ the dual spaces of $H^{k}(G)$, $H^{k}(\Gamma)$, and $\mathbb{H}^{k}$, respectively, for $k=1,2$, where the duality is taken with respect to the pivot spaces $L^{2}(G)$, $L^{2}(\Gamma)$, and $\mathbb{L}^{2}$. For notational convenience, when we write
$(z,z_{\Gamma})=(z_{1},z_{2})\quad\text{in }G\times\Gamma,\quad\text{a.s.},$
we mean that
$z=z_{1}\quad\text{in }G\quad\text{and}\quad z_{\Gamma}=z_{2}\quad\text{on }\Gamma,\quad\text{a.s.}$
Throughout this paper, the boundary vector fields
$\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}:=\left(\sum_{k=1}^{N}a^{jk}_{\Gamma}(x)\,D_{k}z_{\Gamma}\right)_{1\leq j\leq N}\quad\text{and}\quad F_{\Gamma}:=\bigl({F_{\Gamma}}_{j}\bigr)_{1\leq j\leq N},$
@@
appearing in (1.2), are understood as tangential vector fields on the manifold $\Gamma$. More precisely, these vectors are regarded as functions $f_{\Gamma}:\Gamma\to\mathbb{R}^{N}$ such that, for each $x\in\Gamma$,
$f_{\Gamma}(x)\in T_{x}\Gamma:=\bigl\{y\in\mathbb{R}^{N}\,\big|\,y\cdot\nu(x)=0\bigr\}.$ (1.4)
This tangentiality assumption plays a crucial role in the derivation of the Carleman estimate (1.8) for (1.2), as it allows for significant simplifications in the treatment of boundary integral terms. In particular, it permits the use of the standard tangential divergence formula
$\int_{\Gamma}(\nabla_{\Gamma}\cdot f_{\Gamma})\,z_{\Gamma}\,d\sigma=-\int_{\Gamma}f_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma,$
which holds for $f_{\Gamma}\in H^{1}(\Gamma;T_{x}\Gamma)$ and $z_{\Gamma}\in H^{1}(\Gamma)$. In the case where the tangentiality condition (1.4) is not satisfied, we refer to Remark 1.1 for further details.
@@
In Section 2, we prove that equation (1.2) is well posed. More precisely, for any given terminal data $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$ and source terms $F_{1}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$, $F_{2}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, $F_{j}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$, and ${F_{\Gamma}}_{j}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, $j=1,2,\cdots,N$, the system (1.2) admits a unique weak solution
@@
$(z,z_{\Gamma};Z,\widehat{Z})\in\Bigl(L^{2}_{\mathcal{F}}\bigl(\Omega;C([0,T];\mathbb{L}^{2})\bigr)\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})\Bigr)\times L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2}).$